# On the spectral norm of Rademacher matrices

### v1

@@ 1. Introduction and main results
Organization of the paper. In Section 2 we discuss basic tools used in the sequel, including an improvement of the van Handel bound for norms of Gaussian matrices from [12]. In Section 3 we derive a weaker version of Theorem 7 with $\log(d_{A})$ instead of $\log\log(d_{A})$ factors. The last section is devoted to the proof of Theorem 5.
@@ 2. Tools
i) The connected subset $I$ may be chosen by first choosing its spanning tree rooted at $v$ and then labelling the vertices of the tree. The number of unlabelled rooted trees is less than the number of oriented trees with $k$ vertices, i.e., less than the $(k-1)$-th Catalan number $C_{k-1}\leq 4^{k-1}$. The root of the tree is $v$ and the rest of vertices may be labelled in at most $d_{H}^{k-1}$ ways.
@@ 2.1. Van Handel-type bound
Let us first formulate and prove a symmetric variant of Propostion 12.
Proposition 13.
@@
Let $(Y_{1}^{*},\ldots,Y_{n}^{*})$ be a nondecreasing rearrangement of $|Y_{1}|,\ldots,|Y_{n}|$. We set $k=n$ if $b^{2}\leq 1/n$, otherwise we choose $11\leq k\leq n-$ 1 such that $\frac{1}{k+1}<b^{2}\leq\frac{1}{k}$. Then $\mathrm{Log}(nb^{2})\sim\mathrm{Log}(n/k)$ and
$\sup_{s\in B_{2}^{n}\cap bB_{\infty}^{n}}\sum_{i=1}^{n}s_{i}^{2}Y_{i}\leq\frac{1}{k}(Y_{1}^{*}+\ldots+Y_{k}^{*}).$
A standard argument shows that ${\mathbb{E}}Y_{l}^{*}\leq({\mathbb{E}}|Y_{l}^{*}|^{2})^{1/2}\lesssim\max_{i}m_{i}+\mathrm{Log}^{1/2}(n/l)\max_{i}\sigma_{i}$. Thus
${\mathbb{E}}\sup_{s\in B_{2}^{n}\cap bB_{\infty}^{n}}\sum_{i=1}^{n}s_{i}^{2}Y_{i}\lesssim\max_{i}m_{i}+\frac{1}{k}\sum_{l=1}^{k}\sqrt{\mathrm{Log}\bigl{(}\frac{n}{l}\bigr{)}}\max_{i}\sigma_{i}\lesssim\max_{i}m_{i}+\sqrt{\mathrm{Log}\bigl{(}\frac{n}{k}\bigr{)}}\max_{i}\sigma_{i}$
and
@@
Proof of Proposition 13.
[…]
@@
Hence the Slepian-Fernique inequality [10, Theorem 3.15]. yields
@@
Variables $Y_{i}$ are centered Gaussian and (see the proof of Corollary 4.2 in [12]) $({\mathbb{E}}Y_{i}^{2})^{1/2}\leq\|(a_{i,j})_{j}\|_{4}$. Hence Lemma 14 applied with $m_{i}=0$ and $\sigma_{i}=\|(a_{i,j})_{j}\|_{4}$ yields
@@
where $V_{j}=\sqrt{\sum_{i}a_{ij}^{2}g_{i}^{2}}$. The Gaussian concentration [10, Lemma 3.1] yields
@@
so Lemma 14 applied with $Y_{j}=V_{j}$, $m_{j}=\|(a_{i,j})_{i}\|_{2}$ and $\sigma_{j}=\|(a_{i,j})_{i}\|_{\infty}$ yields
@@
Proof of Proposition 12.
We apply Proposition 13 to the symmetric $(n+m)\times(n+m)$ matrix $\tilde{A}=(\tilde{a}_{i,j})$ of the form $\tilde{A}=\left(\begin{array}[]{cc}0&A\\ A^{T}&0\end{array}\right)$. Observe that $\max_{i,j}|\tilde{a}_{i,j}|=\max_{i,j}|a_{i,j}|$,
@@ 3. Bounds up to polylog factors
$I^{\prime}=\{j\in[n]\colon\ \exists_{i\in I}\ (i,j)\in E_{A}\},\quad I^{\prime\prime}=\{i\in[n]\colon\ \exists_{i_{0}\in I,j\in[n]}\ (i_{0},j),(i,j)\in E_{A}\}.$ (10)
@@
By Remark 6 we may and will assume that $a_{i,i}=0$ for all $i$.
@@
$X=X(A):=\max_{1\leq k,l\leq n}X_{k,l}=\max_{\emptyset\neq I,J\subset V}\frac{1}{\sqrt{|I||J|}}\max_{\eta_{i},\eta_{j}^{\prime}=\pm 1}\sum_{i\in I,j\in J}a_{i,j}\varepsilon_{i,j}\eta_{i}\eta_{j}.$ (11)
Lemma 15.
@@
Let us first fix sets $I\in\mathcal{I}(k)$ and $J\in\mathcal{I}(l)$. Let $I_{1},\ldots,I_{r}$ be connected components of $I\cap J^{\prime}$ in $G_{2}$ and $J_{u}:=J\cap I_{u}^{\prime}$. Then sets $J_{1},\ldots,J_{r}$ are disjoint and 4-connected subsets of $J$. Hence, for every $\eta_{i},\eta_{j}^{\prime}=\pm 1$ we have
@@
$X_{k,l}\leq\max_{k^{\prime}\leq k,l^{\prime}\leq l}\overline{X}_{k^{\prime},l^{\prime}}.$ (12)
@@
Obviously $\overline{X}_{k^{\prime},l^{\prime}}\leq\|(a_{i,j}\varepsilon_{i,j})\|$, so by (7) we may assume that $n\geq d_{A}\geq 3$. By the symmetry it is enough to consider only the case $l\geq k$. By Lemma 11, $2^{k}|\mathcal{I}_{4}(k)|\leq n(8d_{A}^{4})^{k}\leq nd_{A}^{6k}$.
@@
The assertion follows by Proposition 10. ∎
Corollary 17.
@@
${\mathbb{E}}X={\mathbb{E}}\max_{\emptyset\neq I,J\subset V}\frac{1}{\sqrt{|I||J|}}\max_{\eta_{i},\eta_{j}^{\prime}=\pm 1}\sum_{i\in I,j\in J}a_{i,j}\varepsilon_{i,j}\eta_{i}\eta_{j}\lesssim\sqrt{\mathrm{Log}\,d_{A}}\max_{i}\|(a_{i,j})_{j}\|_{2}+R_{A}(\mathrm{Log}\,n).$
@@
Lemmas 15 and 16 imply that for a fixed $1\leq k,l\leq n$
@@
and the assertion follows by Proposition 10. ∎
Proposition 18.
@@
By Remark 6 we may assume that $a_{i,i}=0$ for all $i$ and $n\geq d_{A}\geq 3$.
@@
$I_{k}(s)=\{i\in V\colon\ e^{-k-1}<|s_{i}|\leq e^{-k}\},\quad J_{l}(t)=\{j\in V\colon\ e^{-l-1}<|t_{j}|\leq e^{-l}\}.$
@@
$\bigl{\|}(a_{i,j}\varepsilon_{i,j})_{i,j\in V}\bigr{\|}\leq\sup_{\|s\|_{2},\|t\|_{2}\leq 1}\sum_{k,l}e^{-k-l}\sup_{\eta_{i},\eta_{j}^{\prime}=\pm 1}\sum_{i\in I_{k}(s),j\in J_{l}(t)}a_{i,j}\varepsilon_{i,j}\eta_{i}\eta^{\prime}_{j}.$
@@
and the assertion follows by Corollary 17. ∎
4. Proof of Theorem 5
By Remark 6 we may assume that $a_{i,i}=0$ for all $i$ and $n\geq d_{A}\geq 3$.
@@ 4. Proof of Theorem 5
By symmetry it is enough to bound only the terms with $r\geq 0$. Let $X$ be defined by (11).
@@
where the last inequality follows by (13).
Hence Corollary 17 yields
@@
${\mathbb{E}}\sup_{\|s\|_{2}\leq 1}\sup_{\|t\|_{2}\leq 1}\sum_{k\geq 1}\sum_{i\in I_{k}(s)}\sum_{j\in J_{k+r}(t)}\mathbbm{1}_{\{|J_{k+r}(t)|\geq d_{A}^{(2r-5)/40}|I_{k}(s)|\}}$ $a_{i,j}\varepsilon_{i,j}s_{i}t_{j}$
@@
Recall definition (10) of sets $I^{\prime}$ and $I^{\prime\prime}$. Let us fix $0\leq r\leq 40$ and $k,t,s$ such that $d_{A}^{(2r-5)/40}<|J_{k+r}(t)|/|I_{k}(s)|<d_{A}^{(2r+5)/40}$. Let $|I_{k}(s)|=m$. Let us consider the following greedy algorithm with output being a subset $\{i_{1},\ldots,i_{M}\}$ of $I_{k}(s)$ of size $M\leq m$

### v2

@@ 1. Introduction and main results
Organization of the paper. In the next section we present several examples of applications of the main results. In Section 3 we discuss basic tools used in the sequel, including an improvement of the van Handel bound for norms of Gaussian matrices from [17]. In Section 4 we derive a weaker version of Theorem 1.8 with $\log(d_{A})$ instead of $\log\log(d_{A})$ factors. The last section is devoted to the proof of Theorem 1.6.
2. Examples
In the first examples we discuss how to apply Corollary 1.3 to estimate ${\mathbb{E}}\|(\mathbbm{1}_{E}(i,j)\varepsilon_{i,j})\|$ for various classes of graphs $G=([n],E)$. Due to Remark 1.2 the presented bounds are valid also for symmetric random matrices ${\mathbb{E}}\|(\mathbbm{1}_{E}(i,j)\tilde{\varepsilon}_{i,j})_{i,j}\|$.
Example 2.1.
Let $G=([n],E)$ be a graph with maximal degree $d$. Then
(2.1) ${\mathbb{E}}\|(\mathbbm{1}_{E}(i,j)\varepsilon_{i,j})_{i,j\leq n}\|\leq d$
and
(2.2) ${\mathbb{E}}\|(\mathbbm{1}_{E}(i,j)\varepsilon_{i,j})_{i,j\leq n}\|\lesssim\sqrt{d}+\sqrt{\mathrm{Log}\,n}.$
The first bound follows by (1.10). Estimate (2.2) follows from (1.1), one may also use (1.6) and bound the operator norm of $\mathbbm{1}_{F}$ by the Hilbert-Schmidt norm.
Example 2.2.
Let $G=([n],E)$ be the disjoint union of $m$ complete graphs of size $d+1$ (i.e. $n=m(d+1)$ and $\mathbbm{1}_{E}$ is a block-diagonal matrix with $m$ $(d+1)\times(d+1)$ blocks with ones outside diagonal). Then
${\mathbb{E}}\|(\mathbbm{1}_{E}(i,j)\varepsilon_{i,j})\|\sim\sqrt{d}+\min\{d,\sqrt{\mathrm{Log}\,n}\}\sim\begin{cases}d&\mbox{if }d\leq\sqrt{\mathrm{Log}\,n},\\ \sqrt{\mathrm{Log}\,n}&\mbox{if }\sqrt{\mathrm{Log}\,n}\leq d\leq\mathrm{Log}\,n,\\ \sqrt{d}&\mbox{if }d\geq\mathrm{Log}\,n.\end{cases}$
Indeed, $G$ has degree $d$ and
$\max_{F\subset E,|F|\leq\mathrm{Log}\,n}\|(\mathbbm{1}_{\{(i,j)\in F\}})_{i,j}\|=\max_{F\subset[d+1]\times[d+1],|F|\leq\mathrm{Log}\,n}\|(\mathbbm{1}_{\{(i,j)\in F,i\neq j\}})_{i,j}\|\sim\min\{d,\sqrt{\mathrm{Log}\,n}\}.$
hence the upper bound for ${\mathbb{E}}\|(a_{i,j}\varepsilon_{i,j})\|$ follows by (1.6). To derive the lower bound it is enough to consider only the case $\mathrm{Log}\,n\geq d$. We then apply estimate (1.7) and observe that
$\min_{|I|\leq m-1}\max_{F\subset E,|F|\leq\mathrm{Log}(m-1)}\|(\mathbbm{1}_{\{(i,j)\in F,i,j\notin I\}})_{i,j}\|$ $=\max_{F\subset[d]\times[d],|F|\leq\mathrm{Log}(m-1)}\|(\mathbbm{1}_{\{(i,j)\in F,i\neq j\}})_{i,j}\|$
$\sim\min\{d,\sqrt{\mathrm{Log}\,m}\}\sim\min\{d,\sqrt{\mathrm{Log}\,n}\}.$
Example 2.3.
Let $G=([n],E)$ have maximal degree $d$ and girth at least $\varepsilon\mathrm{Log}\mathrm{Log}\,n$. Then ${\mathbb{E}}\|(\mathbbm{1}_{E}(i,j)\varepsilon_{i,j})\|\sim_{\varepsilon}\sqrt{d}$.
By Corollary (1.3) it is enough to show that any subgraph $H=(V,F)$ of $([n],E)$ with at most $\mathrm{Log}\,n$ edges has spectral radius at most $C(\varepsilon)\sqrt{d}$. Subgraph $H$ has at most $2\mathrm{Log}\,n$ vertices, maximal degree at most $d$ and girth at least $\varepsilon\mathrm{Log}\mathrm{Log}\,n$. Let $k=2\lfloor\frac{\varepsilon}{2}\mathrm{Log}\mathrm{Log}\ n\rfloor$ then $\|(\mathbbm{1}_{F})\|\leq(\mathrm{tr}(\mathbbm{1}_{F}^{k}))^{1/k}$. Since $F$ does not contain cycles of lenghts at most $k$, we have
$\mathrm{tr}(\mathbbm{1}_{F}^{k})=\sum_{i_{1},\ldots,i_{k}}\mathbbm{1}_{\{(i_{1},i_{2})\in F\}}\cdots\mathbbm{1}_{\{(i_{k-1},i_{k})\in F\}}\mathbbm{1}_{\{(i_{k},i_{1})\in F\}}\leq 2\mathrm{Log}\,n\cdot N_{k,d},$
where $N_{k,d}\leq 2^{k}d^{k/2}$ is the number of closed pathes from root to itself in the $d$-regular tree. Hence we obtain the desired bound
$\|(\mathbbm{1}_{F})\|\leq(2\mathrm{Log}\,n)^{1/k}N_{k,d}^{1/k}\lesssim_{\varepsilon}\sqrt{d}.$
The next example generalizes the previous one. It is close to [14, Theorem 1.2], where there was a stronger assumption on the neighborhood diameter, but a more precise bound on the operator norm.
Example 2.4.
Let $G=([n],E)$ have maximal degree $d$ and suppose that the $r$-neighborhood of every vertex contains at most one cycle, where $r\geq\varepsilon\mathrm{Log}\mathrm{Log}\,n$. Then ${\mathbb{E}}\|(\mathbbm{1}_{E}(i,j)\varepsilon_{i,j})\|\sim_{\varepsilon}\sqrt{d}$.
Observe that all cycles of lenght at most $2r$ do not intersect. Let us remove one edge from each cycle of length at most $2r$ and let $G_{1}:=([n],E_{1})$, where $E_{1}$ contains all removed edges and $G_{2}:=([n],E_{2})$, where $E_{2}=E\setminus E_{1}$. Then $G_{2}$ has girth at least $2r$ and the maximal degree of $G_{1}$ is at most $1$. Thus by (2.1) and Example 2.3,
${\mathbb{E}}\|(\mathbbm{1}_{E}(i,j)\varepsilon_{i,j})\|\leq{\mathbb{E}}\|(\mathbbm{1}_{E_{1}}(i,j)\varepsilon_{i,j})\|+{\mathbb{E}}\|(\mathbbm{1}_{E_{2}}(i,j)\varepsilon_{i,j})\|\lesssim_{\varepsilon}1+\sqrt{d}\lesssim\sqrt{d}.$
The next example is similar to [14, Theorem 1.16], where it was showned that for a random $d$-regular graph $G=([n],E)$, $\|(\mathbbm{1}_{E}(i,j)\tilde{\varepsilon}_{i,j})\|\leq 2\sqrt{d-1}+\varepsilon$ with probability $1-o_{n}(1)$.
Example 2.5.
Let $G=([n],E)$ be a random $d$-regular graph. Then ${\mathbb{E}}\|(\mathbbm{1}_{E}(i,j)\varepsilon_{i,j})\|\sim\sqrt{d}$.
By (2.2) we may assume that $\mathrm{Log}\,n\geq d$ and by (2.1) that $d\geq 3$. Following [8] we say that a graph is $r$-tangle free if the $r$-neighborhood of every vertex contains at most one cycle By Example 2.4 and (2.1) we have
${\mathbb{E}}\|(\mathbbm{1}_{E}(i,j)\varepsilon_{i,j})\|\leq C\sqrt{d}\,\mathbb{P}(G\mbox{ is $\mathrm{Log}\mathrm{Log}n$-tangle free})+d\,\mathbb{P}(G\mbox{ is not $\mathrm{Log}\mathrm{Log}n$-tangle free}).$
By [8, Lemma 27]
$\mathbb{P}(\mbox{$G$ is not $\mathrm{Log}\mathrm{Log}n$-tangle free})\lesssim\frac{(d-1)^{\mathrm{Log}\mathrm{Log}n}}{n}\lesssim d^{-1/2},$
where the last bound follows since we assume that $\mathrm{Log}\,n\geq d$.
The next example essentially recovers [2, Corollary 1.3].
Example 2.6.
Let $A$ be the adjacency matrix of a $d$-regular graph and suppose that all eigenvalues of $A$ besides the largest one are bounded in absolute value by $\lambda$. Then ${\mathbb{E}}\|(a_{i,j}\varepsilon_{i,j})\|\lesssim\lambda$.
We have $\lambda\gtrsim\sqrt{d}$ by the Alon-Boppana theorem, so by (2.2) we may assume that $\mathrm{Log}\,n\geq d$. By (1.6) it is enough to show that any subgraph $H=(V,F)$ of $([n],E)$ with at most $\mathrm{Log}\,n$ edges has spectral radius at most $C\lambda$. Subgraph $H$ has at most $2\mathrm{Log}\,n$ vertices. Let $w=n^{-1/2}(1,\ldots,1)$ be the eigenvector of $A$ corresponding to the largest eingenvalue $d$. Any $v\in{\mathbb{R}}^{V}$ with $\|v\|_{2}=1$ we may represent as $v=\langle v,w\rangle w+v^{\prime}$, where $\langle v^{\prime},w\rangle=0$ and $\|v^{\prime}\|_{2}\leq 1$. Thus
$\|\mathbbm{1}_{F}v\|_{2}\leq\|\mathbbm{1}_{F}v^{\prime}\|+\langle v,w\rangle\|\mathbbm{1}_{F}w\|_{2}\leq\lambda+\|v\|_{1}n^{-1/2}|V|^{1/2}dn^{-1/2}\leq\lambda+2d\frac{\mathrm{Log}\,n}{n}\lesssim\lambda,$
where the last estimate follows since $\mathrm{Log}\,n\geq d$ and $\lambda\gtrsim 1$.
The next example shows that estimate (1.6) cannot be reversed.
Example 2.7.
Let $1\leq d\leq n$ and $E=([d]\times[d])\cup\{(i,i)\colon\ d<i\leq n\}$ (i.e $\mathbbm{1}_{E}$ is block diagonal with one $d\times d$ block of ones and ndn-d blocks of single ones). Then
$\sup_{F\subset E,|F|\leq\mathrm{Log}\,n}\|(\mathbbm{1}_{\{(i,j)\in F\}})_{i,j}\|=\sup_{F\subset[d]\times[d],|F|\leq\mathrm{Log}\,n}\|(\mathbbm{1}_{\{(i,j)\in F\}})_{i,j}\|\sim\min\{d,\sqrt{\mathrm{Log}\,n}\}$
and the RHS of (1.6) is of the order $\sqrt{d}+\min\{d,\sqrt{\mathrm{Log}\,n}\}$, whereas
${\mathbb{E}}\|(\mathbbm{1}_{E}(i,j)\varepsilon_{i,j})_{i,j\leq n}\|={\mathbb{E}}\|(\varepsilon_{i,j})_{i,j\leq d}\|\sim\sqrt{d}.$
The last example concerns randomized circulant matrices, investigated in [11].
Example 2.8.
Suppose that $(a_{i,j})$ is a circulant matrix, i.e. $a_{i,j}=b_{i-j\,\mathrm{mod}\,n}$ for a deterministic sequence $(b_{i})_{i=0}^{n-1}$. Then for any $i$ and $j$, $\|(a_{i,j})_{j}\|_{2}=\|(a_{i,j})_{i}\|_{2}=\|(b_{i})_{i}\|_{2}$. Moreover, as was shown in the proof of [11, Theorem 1.3]
$\sup_{\|s\|_{2},\|t\|_{2}\leq 1}\left\|\sum_{i,j}a_{ij}\varepsilon_{ij}s_{i}t_{j}\right\|_{\mathrm{Log}\,n}\lesssim\inf_{I\subset[n],|I|\leq n/4}\sup_{\|s\|_{2},\|t\|_{2}\leq 1}\left\|\sum_{i,j\not\in I}a_{ij}\varepsilon_{ij}s_{i}t_{j}\right\|_{\mathrm{Log}(n/4)},$
Thus Theorem 1.9 improves [11, Theorem 1.3] and yields that for circulant matrices
$\|(b_{i})\|_{2}+R_{A}(\mathrm{Log}\,n)\lesssim{\mathbb{E}}\|(a_{i,j}\varepsilon_{i,j})\|\lesssim\mathrm{Log}\mathrm{Log}\mathrm{Log}\,n\big{(}\|(b_{i})\|_{2}+R_{A}(\mathrm{Log}\,n)\big{)}.$
@@ 3. Tools
i) We choose a connected subset $I\ni v$ by constructing its spanning tree, rooted at $v$. In order to do it we first choose an unlabelled rooted tree with $k$ vertices and then label its vertices by elements of $V_{H}$. The number of unlabelled rooted trees is less than the number of ordered trees with $k$ vertices, i.e., less than the (k1)(k-1)-th Catalan number $C_{k-1}\leq 4^{k-1}$. The root of the tree has label $v$ and the rest of vertices may be labelled in at most $d_{H}^{k-1}$ ways.
@@ 3.1. Van Handel-type bound
Let us first formulate and prove a symmetric variant of Propostion $3.$
Proposition 3.4.
@@
Let $(x_{1}^{*},\ldots,x_{n}^{*})$ denote a nondecreasing rearrangement of $|x_{1}|,\ldots,|x_{n}|$. We set $k=n$ if $b^{2}\leq 1/n$, otherwise we choose $1\leq k\leq n-1$ such that $\frac{1}{k+1}<b^{2}\leq\frac{1}{k}$. Then $\mathrm{Log}(nb^{2})\sim\mathrm{Log}(n/k)$ and for any $s\in B_{2}^{n}\cap bB_{\infty}^{n}$,
$\sum_{i=1}^{n}s_{i}^{2}Y_{i}$ $\leq\sum_{i=1}^{n}|s_{i}^{*}|^{2}Y_{i}^{*}\leq\sum_{i=1}^{k}|s_{i}^{*}|^{2}Y_{i}^{*}+\Bigl{(}1-\sum_{i=1}^{k}|s_{i}^{*}|^{2}\Bigr{)}Y_{k}^{*}$
$=\sum_{i=1}^{k}\Bigl{(}|s_{i}^{*}|^{2}Y_{i}^{*}+\Bigl{(}\frac{1}{k}-|s_{i}^{*}|^{2}\Bigr{)}Y_{k}^{*}\Bigr{)}\leq\frac{1}{k}(Y_{1}^{*}+\ldots+Y_{k}^{*}).$
For any $1\leq l\leq n$ and $u\geq 0$,
$\mathbb{P}(Y_{l}^{*}\geq\max_{i}m_{i}+u\max_{i}\sigma_{i})\leq\frac{1}{l}\sum_{i=1}^{n}\mathbb{P}(Y_{i}\geq\max_{i}m_{i}+u\max_{i}\sigma_{i})\leq\frac{n}{l}e^{-u^{2}/2}.$
Hence integration by parts yields ${\mathbb{E}}Y_{l}^{*}\leq({\mathbb{E}}|Y_{l}^{*}|^{2})^{1/2}\lesssim\max_{i}m_{i}+\mathrm{Log}^{1/2}(n/l)\max_{i}\sigma_{i}$. Thus
${\mathbb{E}}\sup_{s\in B_{2}^{n}\cap bB_{\infty}^{n}}\sum_{i=1}^{n}s_{i}^{2}Y_{i}$ $\leq\frac{1}{k}\sum_{l=1}^{k}{\mathbb{E}}Y_{l}^{*}\lesssim\max_{i}m_{i}+\frac{1}{k}\sum_{l=1}^{k}\sqrt{\mathrm{Log}\bigl{(}\frac{n}{l}\bigr{)}}\max_{i}\sigma_{i}$
$\lesssim\max_{i}m_{i}+\sqrt{\mathrm{Log}\bigl{(}\frac{n}{k}\bigr{)}}\max_{i}\sigma_{i}.$
Similarly,
@@
Proof of Proposition 3.4.
Let $(g_{i,j})_{i,j\leq n}$ be a symmetric Gaussian matrix (i.e., $g_{i,j}=g_{j,i}$ and $(g_{i,j})_{i\geq j}$ are iid $\mathcal{N}(0,1)$ r.v’s), independent of $\tilde{\varepsilon}_{i,j}$. We have for any matrix norm $\|\cdot\|$,
@@
Hence the Slepian-Fernique inequality [13, Theorem 3.15]. yields
@@
Variables $Y_{i}$ are centered Gaussian and (see the proof of Corollary 4.2 in [17]) $({\mathbb{E}}Y_{i}^{2})^{1/2}\leq\|(a_{i,j})_{j}\|_{4}$. Hence Lemma 3.5 applied with $m_{i}=0$ and $\sigma_{i}=\|(a_{i,j})_{j}\|_{4}$ yields
@@
where $V_{j}=\sqrt{\sum_{i}a_{ij}^{2}g_{i}^{2}}$. The Gaussian concentration [13, Lemma 3.1] yields
@@
so Lemma 3.5 applied with $Y_{j}=V_{j}$, $m_{j}=\|(a_{i,j})_{i}\|_{2}$ and $\sigma_{j}=\|(a_{i,j})_{i}\|_{\infty}$ yields
@@
Proof of Proposition $3.$
We apply Proposition 3.4 to the symmetric $(n+m)\times(n+m)$ matrix $\tilde{A}=(\tilde{a}_{i,j})$ of the form $\tilde{A}=\left(\begin{array}[]{cc}0&A\\ A^{T}&0\end{array}\right)$. Observe that $\max_{i,j}|\tilde{a}_{i,j}|=\max_{i,j}|a_{i,j}|$,
@@ 4. Bounds up to polylog factors
In this section we derive weaker estimates than in Theorem 1.8 (with powers of $\log d_{A}$ instead of $\log\log(d_{A})$). They will be used in the proof of Theorem 1.6 to estimate the parts of Bernoulli process $(\sum_{i,j}a_{i,j}\varepsilon_{i,j}s_{i}t_{j})_{s,t\in B_{2}^{n}}$, where coefficients $s_{i}$ and $t_{j}$ are of the same order.
@@
(4.1) $I^{\prime}=\{j\in[n]\colon\ \exists_{i\in I}\ (i,j)\in E_{A}\},\quad I^{\prime\prime}=\{i\in[n]\colon\ \exists_{i_{0}\in I,j\in[n]}\ (i_{0},j),(i,j)\in E_{A}\}.$
@@
By Remark 1.7 we may and will assume that $a_{i,i}=0$ for all $i$.
@@
(4.2) $X=X(A):=\max_{1\leq k,l\leq n}X_{k,l}=\max_{\emptyset\neq I,J\subset V}\frac{1}{\sqrt{|I||J|}}\max_{\eta_{i},\eta^{\prime}_{j}=\pm 1}\sum_{i\in I,j\in J}a_{i,j}\varepsilon_{i,j}\eta_{i}\eta^{\prime}_{j}.$
Observe that $X$ is nonnegative.
Variables $\overline{X}_{k,l}$ are easier to estimate than $X_{k,l}$, since the number of $4$-connected subsets is much smaller than the number of all subsets – calculations based on this idea are made in Lemma 4.2. Lemma 4.1 shows that in expectation these variables do not differ too much.
Lemma 4.1.
@@
Let us first fix sets $I\in\mathcal{I}(k)$ and $J\in\mathcal{I}(l)$. Let $I_{1},\ldots,I_{r}$ be connected components of $I\cap J^{\prime}$ in $G_{2}$ and $J_{u}:=J\cap I_{u}^{\prime}$. Then sets $J_{1},\ldots,J_{r}$ are disjoint. They are also 4-connected subsets of $J$, since otherwise there would exists a nonempty set $V\subsetneq J_{u}$ such that $\rho_{A}(V,J_{u}\setminus V)\geq 4$ and taking for $\tilde{V}$ neighbours of $V$ in $I_{u}$ we would have $\emptyset\neq\tilde{V}\subsetneq I_{u}$ and $\rho_{A}(\tilde{V},I_{u}\setminus\tilde{V})\geq 2$, contradicting 2-connectivity of $I_{u}$. Hence, for every $\eta_{i},\eta_{j}^{\prime}=\pm 1$ we have
@@
(4.3) $X_{k,l}\leq\max_{k^{\prime}\leq k,l^{\prime}\leq l}\overline{X}_{k^{\prime},l^{\prime}}.$
@@
Obviously $\overline{X}_{k^{\prime},l^{\prime}}\leq\|(a_{i,j}\varepsilon_{i,j})\|$, so by (1.10) we may assume that $n\geq d_{A}\geq 3$. By the symmetry it is enough to consider only the case $l\geq k$. By Lemma 3.2, $2^{k}|\mathcal{I}_{4}(k)|\leq n(8d_{A}^{4})^{k}\leq nd_{A}^{6k}$.
@@
The assertion follows by Proposition 3.1. ∎
Corollary 4.3.
@@
${\mathbb{E}}X={\mathbb{E}}\max_{\emptyset\neq I,J\subset V}\frac{1}{\sqrt{|I||J|}}\max_{\eta_{i},\eta_{j}^{\prime}=\pm 1}\sum_{i\in I,j\in J}a_{i,j}\varepsilon_{i,j}\eta_{i}\eta^{\prime}_{j}\lesssim\sqrt{\mathrm{Log}\,d_{A}}\max_{i}\|(a_{i,j})_{j}\|_{2}+R_{A}(\mathrm{Log}\,n).$
@@
Lemmas 4.1 and 4.2 imply that for a fixed $1\leq k,l\leq n$
@@
and the assertion follows by Proposition 3.1. ∎
Proposition 4.4.
@@
By Remark 1.7 we may assume that $a_{i,i}=0$ for all $i$ and $n\geq d_{A}\geq 3$.
@@
$I_{k}(s)=\{i\leq n\colon\ e^{-k-1}<|s_{i}|\leq e^{-k}\},\quad J_{l}(t)=\{j\leq n\colon\ e^{-l-1}<|t_{j}|\leq e^{-l}\}.$
@@
$\bigl{\|}(a_{i,j}\varepsilon_{i,j})_{i,j\leq n}\bigr{\|}\leq\sup_{\|s\|_{2},\|t\|_{2}\leq 1}\sum_{k,l}e^{-k-l}\sup_{\eta_{i},\eta_{j}^{\prime}=\pm 1}\sum_{i\in I_{k}(s),j\in J_{l}(t)}a_{i,j}\varepsilon_{i,j}\eta_{i}\eta^{\prime}_{j}.$
@@
and the assertion follows by Corollary 4.3. ∎
5. Proof of Theorem 1.6
By Remark 1.7 we may assume that $a_{i,i}=0$ for all $i$ and $n\geq d_{A}\geq 3$.
@@
By symmetry it is enough to bound only the terms with $r\geq 0$. Let $X$ be defined by (4.2).
@@
where the last inequality follows by (5.1).
Hence Corollary 4.3 yields
@@
${\mathbb{E}}\sup_{\|s\|_{2}\leq 1}\sup_{\|t\|_{2}\leq 1}\sum_{k\geq 1}\sum_{i\in I_{k}(s)}\sum_{j\in J_{k+r}(t)}\mathbbm{1}_{\{|J_{k+r}(t)|\leq d_{A}^{(2r-5)/40}|I_{k}(s)|\}}$ $a_{i,j}\varepsilon_{i,j}s_{i}t_{j}$
@@
Recall definition (4.1) of sets $I^{\prime}$ and $I^{\prime\prime}$. Let us fix $0\leq r\leq 40$ and $k,t,s$ such that $d_{A}^{(2r-5)/40}<|J_{k+r}(t)|/|I_{k}(s)|<d_{A}^{(2r+5)/40}$. Let $|I_{k}(s)|=m$. Let us consider the following greedy algorithm with output being a subset $\{i_{1},\ldots,i_{M}\}$ of $I_{k}(s)$ of size $M\leq m$